# Shuffle Products, Degenerate Affine Hecke Algebras, and Quantum Toda Lattice

### v1

@@ 5.2. Other representations
In this subsection, we compare the isomorphism of Proposition 5.7 with the trivialization (5.3) provided by the Kirillov projector.
@@
In what follows, we will drop the tensor product sign by considering $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes(\mathbf{C}^{N})^{\otimes k}$ as a submodule of $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\ltimes\bigoplus_{i=0}^{\infty}(\mathbf{C}^{N})^{\otimes i}$ and using notations of 5.6.
@@
Take $1\leq\alpha\leq k$ such that $i_{\alpha-1}\geq i_{\alpha}+1\geq i_{\alpha+1}$ (i.e. the non-increasing condition is still satisfied). By 5.4, we have
@@
due to the properties of Theorem 5.2 of the Kirillov projector. Consider the sum for $\beta\leq i_{\alpha}$. Since we assumed that $i_{1}\geq\ldots\geq i_{\alpha-1}>i_{\alpha}$, the elements $E_{i_{\alpha}\beta}$ also commute with the tail $v_{i_{1}}\ldots v_{i_{\alpha-1}}$. Using again the properties of the Kirillov projector, we have
@@ 5.3. Mirabolic reduction and Feigin-Odesskii shuffle algebras
a mirabolic version of the Kostant-Whittaker reduction from 5.1 for free Harish-Chandra bimodules. It follows from [KalmykovYangians, Theorem 5.3] that the action of the Kirillov projector of Theorem 5.2 is well-defined on $\mathrm{U}_{\hbar}(\mathfrak{m}_{N})\otimes V/\mathfrak{n}_{+}^{\psi}$ and gives an isomorphism
@@
The following is an immediate corollary of 5.4 and Proposition 5.10.
@@
Therefore, by Proposition 5.11 and Proposition 2.7, we have
@@
Therefore, by Proposition 5.11 and Proposition 2.7, we have
@@
where $s_{\lambda}(\vec{\Omega})$ is the Schur polynomial associated to partition $\lambda$, see [Macdonald, Section I.3], and the factor can be viewed as a deformation of the Vandermonde determinant.
By [KalmykovYangians, Theorem 6.3], the functor (5.9) is monoidal, in particular, if $A$ is an algebra in $\mathrm{Rep}(\mathrm{GL}_{N})$, one can define a mirabolic analog $\mathrm{U}_{\hbar}(\mathfrak{m}_{N})\ltimes A$ of 3.4, and its reduction $\mathrm{U}_{\hbar}(\mathfrak{m}_{N})\ltimes A/\!\!/_{\psi}\,N_{+}$ is an algebra as well, similarly to 5.6. Recall the rational Feigin-Odesskii shuffle algebra ${}^{N}\mathcal{S}^{\hbar}$ of 2.9. The following is a corollary of Theorem 2.11, Theorem 5.9, and Example 5.12.
@@
As we mentioned in the introduction, the shuffle algebra ${}^{N}\mathcal{S}^{\hbar}$ is a rational degeneration of its elliptic version, in view of Section 2 and the results of [EndelmanHodgesTwistedSU]. The original construction of [FeiginOdesskiiSklyanin] of the latter used Belavin’s $R$-matrix that can be associated to a stable bundle of rank $N$ and degree 1 (for instance, see “Geometric interpretation” after [LOZ, (1.56)] in loc. cit. for precise statement). Unfortunately, we do not know an interpretation of Theorem 5.14 in these terms. However, it seems to be a step in the right direction based on the following observation.
Let $E$ be an elliptic curve with the origin $0\in E$. Denote by $\xi_{n,r}$ a unique stable bundle $E$ of rank $r$ whose determinant is $\mathcal{O}_{E}(r\cdot 0)$. As it is argued in [FeiginOdesskii], the elliptic shuffle algebra quantizes a certain Poisson bracket on the space of extensions $\mathrm{Ext}^{1}(\xi_{N,1},\xi_{0,1})$. Its ring of functions is the symmetric algebra on
@@
where $\mathcal{O}_{0}$ is the skyscraper sheaf at the origin. Therefore, it transforms $\mathrm{Hom}^{0}(\xi_{-N,1},\xi_{0,1})$ to the space of morphisms $\xi_{1,N}\rightarrow\mathcal{O}_{0}$ which can interpreted as a mirabolic reduction of the fiber of $\xi_{1,N}$ at the origin. To relate to the Kostant-Whittaker reduction, recall that it quantizes the Kostant slice [Kostant]. The latter can be identified with $\mathfrak{h}_{N}^{W}$. Its elliptic version is the coarse moduli space of degree 0 semistable bundles of rank $N$. At the same time, we can complete a non-zero $\varphi\in\mathrm{Hom}^{0}(\xi_{-N,1},\xi_{0,1})$ to an exact triangle
@@
at $0\in E$. Therefore, it seems that $\mathrm{U}_{\hbar}(\mathfrak{m}_{N})\otimes S^{\bullet}\mathbf{C}^{N}/\!\!/_{\psi}\,N_{+}$ should be a certain quantum (and rational) version of the algebra of Hecke modifications of special kind, if it makes sense.
@@ 6. Finite Miura transform
In this section, we relate the Kostant-Whittaker reduction of Section 5 to the parabolic restriction of Section 4 following the construction of [GinzburgKazhdan, Section 6].
@@
Let $X\in\mathrm{HC}_{\hbar}(\mathrm{GL}_{N})$. The Miura bimodule associated to $X$ is the double quotient $\mathfrak{n}_{-}\backslash X^{\mathrm{gen}}/\mathfrak{n}^{\psi}_{+}$.
@@
Recall the Harish-Chandra map of Example 4.4:
@@
Following [GinzburgKazhdan, Section 6], consider the map
(6.2) $\mathrm{U}_{\hbar}(\mathfrak{h}_{N})^{\mathrm{gen}}\otimes_{\mathbf{Z}_{\mathfrak{gl}_{N}}}(X/\!\!/_{\psi}\,N_{+})\xrightarrow{1\otimes p_{X}}\mathrm{U}_{\hbar}(\mathfrak{h}_{N})^{\mathrm{gen}}\otimes_{\mathbf{Z}_{\mathfrak{gl}_{N}}}\mathfrak{n}_{-}\backslash X/\mathfrak{n}^{\psi}_{+}\xrightarrow{\mathrm{act}}\mathfrak{n}_{-}\backslash X^{\mathrm{gen}}/\mathfrak{n}_{+},$
@@
$\mathrm{act}\colon\mathrm{U}_{\hbar}(\mathfrak{h}_{N})^{\mathrm{gen}}\otimes\mathfrak{n}_{-}\backslash X\rightarrow\mathfrak{n}_{-}\backslash X^{\mathrm{gen}}$
@@
For free Harish-Chandra bimodules, the isomorphism of the second part holds without genericity assumption by [GinzburgKazhdan, Theorem 7.1.8].
@@
In view of the map (6.3), the Miura transform induces a $W$-action on $X^{\mathrm{gen}}/\!\!/N_{-}$ extending the one of Example 4.4 such that $X/\!\!/_{\psi}\,N_{+}\subset(X^{\mathrm{gen}}/\!\!/N_{-})^{W}$.
@@
$X/\!\!/_{\psi}\,N_{+}\xrightarrow{[1_{X}]\otimes\mathrm{id}}(\mathfrak{n}_{-}\backslash X^{\mathrm{gen}}/\mathfrak{n}^{\psi}_{+})\otimes(X/\!\!/_{\psi}\,N_{+})\rightarrow(\mathfrak{n}_{-}\backslash X^{\mathrm{gen}}/\mathfrak{n}^{\psi}_{+})$
on the class of identity $1_{X}\in X$ in $\mathfrak{n}_{-}\backslash X^{\mathrm{gen}}/\mathfrak{n}^{\psi}_{+}$, and similarly for (6.3). Then the lemma follows from a simple general result: let $A,B$ be two algebras and $M$ be an $(A,B)$-bimodule with an element $m\in M$ such that the composition
@@
Consider $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\ltimes S^{\bullet}\mathbf{C}^{N}$. Combining Theorem 5.9, Theorem 4.11, and 6.6, we obtain an algebra homomorphism
@@
The map (6.4) coincides with the GKLO map of Theorem 2.16.
@@
It follows from Example 4.4 that
@@
By arguments of [KalmykovYangians, Corollary 7.12] adapted to the positive case, we have
@@ 7. Toda lattice
In this section, we show an isomorphism between the quantum Toda lattice for $\mathrm{GL}_{N}$ and the truncated shifted Yangian from Section 2. The main object of this section is the algebra $\mathrm{D}_{\hbar}(\mathrm{GL}_{N})$ of differential operators from Section 1.3. Recall that it has two $\mathrm{GL}_{N}$ actions.
@@
the corresponding free Harish-Chandra bimodules of Section 3.
@@
the Kostant-Whittaker reduction of Section 5 for left Harish-Chandra bimodules. In particular,
@@
correspondingly the image of $x$ under the projection map $X\rightarrow X/\mathfrak{n}_{+}^{L,-\psi}$ and the action of the Kirillov projector from Theorem 5.2 on the latter. Recall that if $xP_{-\psi}^{L}$ is $N^{L}_{+}$-invariant, then we have $P_{-\psi}^{L}xP_{-\psi}^{L}=xP_{-\psi}^{L}$.
@@
the parabolic restriction of Section 4 and the Kostant-Whittaker reduction respectively for right Harish-Chandra bimodules. In particular, both
@@
Under the embeddings of Example 3.5, we have
@@
Since $v_{1}\in\mathbf{C}^{N}$ (resp. $\phi_{N}\in(\mathbf{C}^{N})^{*}$) is a highest-weight vector, the embeddings of Example 3.5 factor through the maps of $\mathrm{GL}_{N}^{R}$ Harish-Chandra bimodules
@@
The maps (7.1) descend to homomorphisms between corresponding reductions of Section 4 and Section 5. Denote by
@@
Using (7.1) and (7.2), it is enough to prove that $\bar{\phi}_{i}\bar{v}_{j}=\bar{v}_{j}\bar{\phi}_{i}$. This is the content of Proposition A.2. ∎
@@
We prove only the first equality. Recall the quantum comatrix $\hat{T}^{L}(u)$ from Section 1.3 associated to vector fields generating left translations. Since they commute with the ones generating right translations, its matrix entries descend to elements of $\mathrm{D}_{\hbar}(\mathrm{GL}_{N})/\!\!/N_{-}^{R}$, in other words, we have equality
@@
Using Proposition 1.4 and Proposition A.3, we get
@@
In particular, we have $Z_{1a}^{*}Z_{aN}=P^{R}\hat{T}^{L}_{1N}(w_{a}^{R})P_{-\psi}^{L}$. At the same time, one can show analogously to [KalmykovYangians, Proposition 4.8] that
@@
The center $\mathbf{Z}_{\mathfrak{gl}_{N}}$ is embedded in $\mathrm{D}_{\hbar}(\mathrm{GL}_{N})$ as a subalgebra of bi-invariant differential operators, in particular, descends to the quantum Hamiltonian reduction $N^{L}_{+}\,\backslash\!\!\backslash_{-\psi}\,\mathrm{D}_{\hbar}(\mathrm{GL}_{N})/\!\!/_{\psi}\,N_{+}^{R}$. For any matrix $M$, we denote by $[M]_{ij}$ its $(i,j)$-matrix entry.
@@
    The algebra $N^{L}_{+}\,\backslash\!\!\backslash_{-\psi}\,\mathrm{D}_{\hbar}(\mathrm{GL}_{N})/\!\!/_{\psi}\,N_{+}^{R}$ is generated over $\mathbf{Z}_{\mathfrak{gl}_{N}}$ by $[(E^{R})^{k}X]_{1N}P_{\psi}^{R}$ and $[X^{-1}(E^{R})^{k}]_{1N}P_{\psi}^{R}$ for every $10\leq k\leq N-$ 1.
@@
In particular, we see that $N^{L}_{+}\,\backslash\!\!\backslash_{-\psi}\,\mathrm{D}_{\hbar}(\mathrm{GL}_{N})$ is generated over $\mathbf{Z}_{\mathfrak{gl}_{N}}$ by $P_{-\psi}^{L}X_{ij}P_{-\psi}^{L}$ and $P_{-\psi}^{L}(X^{-1})_{ij}P_{-\psi}^{L}$ for all $i,j$. By Proposition 5.7 and its dual version of Proposition B.2, it is equivalently generated by the classes of matrix entries $[X(E^{L})^{k}]_{\alpha N}P_{-\psi}^{L}$ and $[(E^{L})^{k}X^{-1}]_{1\alpha}P_{-\psi}^{L}$ in view of Proposition 7.2, for every $\alpha$ and $k$. Therefore, by (1.14), $N^{L}_{+}\,\backslash\!\!\backslash_{-\psi}\,\mathrm{D}_{\hbar}(\mathrm{GL}_{N})$ is generated by matrix entries $[(E^{R})^{k}X]_{\alpha N}$ and $[X^{-1}(E^{R})^{k}]_{1\alpha}$; for readability, we drop the sign $P_{-\psi}^{L}$.
@@
Therefore, by 5.4, its dual version of Proposition B.1, and Proposition 3.7, the double Hamiltonian reduction is generated by the $N_{+}$-invariant classes
@@
By 4.6 and its dual version (B.6), it is equivalently generated by (7.2), as required. ∎
Recall the algebra of difference operators ${}^{N}\tilde{\mathcal{A}}_{\hbar}$ from Section 4.4. We will need the following result which is an analog of simplicity of Weyl algebras.
@@
$[\mathbf{u}_{i},c(w_{1},\ldots,w_{N})]\mathbf{u}_{i}^{-1}=c(w_{1},\ldots,w_{i}+\hbar,\ldots,w_{N})-c(w_{1},\ldots,w_{i},\ldots,w_{N})$
@@
It is an algebra homomorphism by 7.3 and Proposition 7.4. Surjectivity follows from the second part of 7.5. If it had a kernel, then it would be generated by some ideal in $\mathbf{C}[\hbar]$ by 7.6, in particular, the action of $\mathbf{C}[\hbar]$ on the class of $[1]\in N^{L}_{+}\,\backslash\!\!\backslash_{-\psi}\,\mathrm{D}_{\hbar}(\mathrm{GL}_{N})/\!\!/N_{-}^{R}$ would have a kernel as well. Consider the Bruhat cell $U=B_{+}N_{-}\subset\mathrm{GL}_{N}$, where $B_{+}$ is the Borel subgroup. Since $(N_{+}^{L},N^{R}_{-})$-action is free, we have
@@
Finally, recall the finite Miura transform of 6.4:
@@
Since $\mathrm{D}_{\hbar}(\mathrm{GL}_{N})\cong\mathrm{U}_{\hbar}(\mathfrak{gl}_{N}^{R})\otimes\mathcal{O}(\mathrm{GL}_{N})$ is a free module, it follows from trivializations (4.8) and (5.3) that this map is injective. By exactness of the Kostant-Whittaker reduction [BezrukavnikovFinkelberg, Lemma 4(a)], the induced map
@@
The following theorem is an explicit presentation of the isomorphism from [BravermanFinkelbergNakajima, Appendix B].
@@
[…]
@@
where the first map is (6.4). By naturality of the Miura transform and Theorem 5.9, the square
[…]
@@
[…]
by Theorem B.3. To check the remaining relation (2.9), we can use 7.8 and check it inside
@@
which is a straightforward calculation, for instance, see [BravermanFinkelbergNakajima, Appendix B(v)]. Therefore, we have a map
@@
whose composition with the Miura transform (7.3) coincides with the GKLO map. It is surjective by the first part of 7.5. By 7.8, the target is identified with its image in ${}^{N}\tilde{\mathcal{A}}_{\hbar}$. Then 2.17 gives an isomorphism
@@
In view of Remark 6.5, the $W$-action on $N^{L}_{+}\,\backslash\!\!\backslash_{-\psi}\,\mathrm{D}_{\hbar}(\mathrm{GL}_{N})/\!\!/N_{-}\cong{}^{N}\tilde{\mathcal{A}}_{\hbar}$ is

### v2

@@ 5.2. Other representations
In this subsection, we compare the isomorphism of Proposition˜5.7 with the trivialization (5.3) provided by the Kirillov projector.
@@
In what follows, we will drop the tensor product sign by considering $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes(\mathbf{C}^{N})^{\otimes k}$ as a submodule of $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\ltimes\bigoplus_{i=0}^{\infty}(\mathbf{C}^{N})^{\otimes i}$ and using notations of ˜5.6.
@@
Take $1\leq\alpha\leq k$ such that $i_{\alpha-1}\geq i_{\alpha}+1\geq i_{\alpha+1}$ (i.e. the non-increasing condition is still satisfied). By ˜5.4, we have
@@
due to the properties of Theorem˜5.2 of the Kirillov projector. Consider the sum for $\beta\leq i_{\alpha}$. Since we assumed that $i_{1}\geq\ldots\geq i_{\alpha-1}>i_{\alpha}$, the elements $E_{i_{\alpha}\beta}$ also commute with the tail $v_{i_{1}}\ldots v_{i_{\alpha-1}}$. Using again the properties of the Kirillov projector, we have
@@ 5.3. Mirabolic reduction and Feigin-Odesskii shuffle algebras
a mirabolic version of the Kostant-Whittaker reduction from ˜5.1 for free Harish-Chandra bimodules. It follows from [Kal25, Theorem 5.3] that the action of the Kirillov projector of Theorem˜5.2 is well-defined on $\mathrm{U}_{\hbar}(\mathfrak{m}_{N})\otimes V/\mathfrak{n}_{+}^{\psi}$ and gives an isomorphism
@@
The following is an immediate corollary of ˜5.4 and Proposition˜5.10.
@@
Therefore, by Proposition˜5.11 and Proposition˜2.7, we have
@@
Therefore, by Proposition˜5.11 and Proposition˜2.7, we have
@@
where $s_{\lambda}(\vec{\Omega})$ is the Schur polynomial associated to partition $\lambda$, see [Mac95, Section I.3], and the factor can be viewed as a deformation of the Vandermonde determinant.
By [Kal25, Theorem 6.3], the functor (5.9) is monoidal, in particular, if $A$ is an algebra in $\mathrm{Rep}(\mathrm{GL}_{N})$, one can define a mirabolic analog $\mathrm{U}_{\hbar}(\mathfrak{m}_{N})\ltimes A$ of ˜3.4, and its reduction $\mathrm{U}_{\hbar}(\mathfrak{m}_{N})\ltimes A/\!\!/_{\psi}\,N_{+}$ is an algebra as well, similarly to ˜5.6. Recall the rational Feigin-Odesskii shuffle algebra $^{N}\mathcal{S}^{\hbar}$ of ˜2.9. The following is a corollary of Theorem˜2.11, Theorem˜5.9, and Example˜5.12.
@@
As we mentioned in the introduction, the shuffle algebra $^{N}\mathcal{S}^{\hbar}$ is a rational degeneration of its elliptic version, in view of Section˜2 and the results of [EH04]. The original construction of [OF89] of the latter used Belavin’s $R$-matrix that can be associated to a stable bundle of rank $N$ and degree 1 (for instance, see “Geometric interpretation” after [LOZ15, (1.56)] in loc. cit. for precise statement). Unfortunately, we do not know an interpretation of Theorem˜5.14 in these terms. However, it seems to be a step in the right direction based on the following observation.
Let $E$ be an elliptic curve with the origin $0\in E$. Denote by $\xi_{n,r}$ a unique stable bundle $E$ of rank $r$ whose determinant is $\mathcal{O}_{E}(r\cdot 0)$. As it is argued in [FO98], the elliptic shuffle algebra quantizes a certain Poisson bracket on the space of extensions $\mathrm{Ext}^{1}(\xi_{N,1},\xi_{0,1})$. Its ring of functions is the symmetric algebra on
@@
where $\mathcal{O}_{0}$ is the skyscraper sheaf at the origin. Therefore, it transforms $\mathrm{Hom}^{0}(\xi_{-N,1},\xi_{0,1})$ to the space of morphisms $\xi_{1,N}\rightarrow\mathcal{O}_{0}$ which can be interpreted as a mirabolic reduction of the fiber of $\xi_{1,N}$ at the origin. To relate to the Kostant-Whittaker reduction, recall that it quantizes the Kostant slice [Kos78]. The latter can be identified with $\mathfrak{h}_{N}/W$. Its elliptic version is the coarse moduli space of degree 0 semistable bundles of rank $N$. At the same time, we can complete a non-zero $\varphi\in\mathrm{Hom}^{0}(\xi_{-N,1},\xi_{0,1})$ to an exact triangle
@@
at $0\in E$. Therefore, it seems that $\mathrm{U}_{\hbar}(\mathfrak{m}_{N})\otimes S^{\bullet}\mathbf{C}^{N}/\!\!/_{\psi}\,N_{+}$ should be a certain quantum (and rational) version of the algebra of Hecke modifications of special kind.
@@ 6. Finite Miura transform
In this section, we relate the Kostant-Whittaker reduction of Section˜5 to the parabolic restriction of Section˜4 following the construction of [GK22, Section 6].
@@
[…]
@@
Recall the Harish-Chandra map of Example˜4.4:
@@
Following [GK22, Section 6], consider the map
[…]
@@
[…]
@@
For free Harish-Chandra bimodules, the isomorphism of the second part holds without genericity assumption by [GK22, Theorem 7.1.8].
@@
In view of the map (6.3), the Miura transform induces a $W$-action on $X^{\mathrm{gen}}/\!\!/N_{-}$ extending the one of Example˜4.4 such that $X/\!\!/_{\psi}\,N_{+}\subset(X^{\mathrm{gen}}/\!\!/N_{-})^{W}$.
@@
[…]
@@
Consider $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\ltimes S^{\bullet}\mathbf{C}^{N}$. Combining Theorem˜5.9, Theorem˜4.11, and ˜6.6, we obtain an algebra homomorphism
@@
The map (6.4) coincides with the GKLO map of Theorem˜2.16.
@@
It follows from Example˜4.4 that
@@
By arguments of [Kal25, Corollary 7.12] adapted to the positive case, we have
@@ 7. Toda lattice
In this section, we show an isomorphism between the quantum Toda lattice for $\mathrm{GL}_{N}$ and the truncated shifted Yangian from Section˜2. The main object of this section is the algebra $\mathrm{D}_{\hbar}(\mathrm{GL}_{N})$ of differential operators from Section˜1.3. Recall that it has two $\mathrm{GL}_{N}$ actions.
@@
the corresponding free Harish-Chandra bimodules of Section˜3.
@@
the Kostant-Whittaker reduction of Section˜5 for left Harish-Chandra bimodules. In particular,
@@
correspondingly the image of $x$ under the projection map $X\rightarrow X/\mathfrak{n}_{+}^{L,-\psi}$ and the action of the Kirillov projector from Theorem˜5.2 on the latter. Recall that if $xP_{-\psi}^{L}$ is $N^{L}_{+}$-invariant, then we have $P_{-\psi}^{L}xP_{-\psi}^{L}=xP_{-\psi}^{L}$.
@@
the parabolic restriction of Section˜4 and the Kostant-Whittaker reduction respectively for right Harish-Chandra bimodules. In particular, both
@@
Under the embeddings of Example˜3.5, we have
@@
Since $v_{1}\in\mathbf{C}^{N}$ (resp. $\phi_{N}\in(\mathbf{C}^{N})^{*}$) is a highest-weight vector, the embeddings of Example˜3.5 factor through the maps of $\mathrm{GL}_{N}^{R}$ Harish-Chandra bimodules
@@
The maps (7.1) descend to homomorphisms between corresponding reductions of Section˜4 and Section˜5. Denote by
@@
Using (7.1) and (7.2), it is enough to prove that $\bar{\phi}_{i}\bar{v}_{j}=\bar{v}_{j}\bar{\phi}_{i}$. This is the content of Proposition˜A.3. ∎
@@
We prove only the first equality. Recall the quantum comatrix $\hat{T}^{L}(u)$ from Section˜1.3 associated to vector fields generating left translations. Since they commute with the ones generating right translations, its matrix entries descend to elements of $\mathrm{D}_{\hbar}(\mathrm{GL}_{N})/\!\!/N_{-}^{R}$, in other words, we have equality
@@
Using Proposition˜1.4 and Proposition˜A.4, we get
@@
In particular, we have $Z_{1a}^{*}Z_{aN}=P^{R}\hat{T}^{L}_{1N}(w_{a}^{R})P_{-\psi}^{L}$. At the same time, one can show analogously to [Kal25, Proposition 4.8] that
@@
The center $\mathbf{Z}_{\mathfrak{gl}_{N}}$ is embedded in $\mathrm{D}_{\hbar}(\mathrm{GL}_{N})$ as a subalgebra of bi-invariant differential operators, in particular, descends to the quantum Hamiltonian reduction $N^{L}_{+}\,\backslash\!\!\backslash_{-\psi}\,\mathrm{D}_{\hbar}(\mathrm{GL}_{N})/\!\!/_{\psi}\,N_{+}^{R}$. For any matrix $M$, we denote by $M]_{ij}$ its $(i,j)$-matrix entry.
@@
    The algebra $N^{L}_{+}\,\backslash\!\!\backslash_{-\psi}\,\mathrm{D}_{\hbar}(\mathrm{GL}_{N})/\!\!/_{\psi}\,N_{+}^{R}$ is generated over $\mathbf{Z}_{\mathfrak{gl}_{N}}$ by $(E^{R})^{k}X]_{1N}P_{\psi}^{R}$ and $X^{-1}(E^{R})^{k}]_{1N}P_{\psi}^{R}$ for every $0\leq k\leq N-1$.
@@
In particular, we see that $N^{L}_{+}\,\backslash\!\!\backslash_{-\psi}\,\mathrm{D}_{\hbar}(\mathrm{GL}_{N})$ is generated over $\mathbf{Z}_{\mathfrak{gl}_{N}}$ by $P_{-\psi}^{L}X_{ij}P_{-\psi}^{L}$ and $P_{-\psi}^{L}(X^{-1})_{ij}P_{-\psi}^{L}$ for all $i,j$. By Proposition˜5.7 and its dual version of Proposition˜B.2, it is equivalently generated by the classes of matrix entries $X(E^{L})^{k}]_{\alpha N}P_{-\psi}^{L}$ and $(E^{L})^{k}X^{-1}]_{1\alpha}P_{-\psi}^{L}$ in view of Proposition˜7.2, for every $\alpha$ and $k$. Therefore, by (1.14), $N^{L}_{+}\,\backslash\!\!\backslash_{-\psi}\,\mathrm{D}_{\hbar}(\mathrm{GL}_{N})$ is generated by matrix entries $(E^{R})^{k}X]_{\alpha N}$ and $X^{-1}(E^{R})^{k}]_{1\alpha}$; for readability, we drop the sign $P_{-\psi}^{L}$.
@@
Therefore, by ˜5.4, its dual version of Proposition˜B.1, and Proposition˜3.7, the double Hamiltonian reduction is generated by the $N_{+}$-invariant classes
@@
By ˜4.6 and its dual version (B.6), it is equivalently generated by (7.2), as required. ∎
Recall the algebra of difference operators $^{N}\tilde{\mathcal{A}}_{\hbar}$ from Section˜4.4. We will need the following result which is an analog of simplicity of Weyl algebras.
@@
$\mathbf{u}_{i},c(w_{1},\ldots,w_{N})]\mathbf{u}_{i}^{-1}=c(w_{1},\ldots,w_{i}+\hbar,\ldots,w_{N})-c(w_{1},\ldots,w_{i},\ldots,w_{N})$
@@
It is an algebra homomorphism by ˜7.3 and Proposition˜7.4. Surjectivity follows from the second part of ˜7.5. If it had a kernel, then it would be generated by some ideal in $\mathbf{C}[\hbar]$ by ˜7.6, in particular, the action of $\mathbf{C}[\hbar]$ on the class of $1]\in N^{L}_{+}\,\backslash\!\!\backslash_{-\psi}\,\mathrm{D}_{\hbar}(\mathrm{GL}_{N})/\!\!/N_{-}^{R}$ would have a kernel as well. Consider the Bruhat cell $U=B_{+}N_{-}\subset\mathrm{GL}_{N}$, where $B_{+}$ is the Borel subgroup. Since $(N_{+}^{L},N^{R}_{-})$-action is free, we have
@@
Finally, recall the finite Miura transform of ˜6.4:
@@
Since $\mathrm{D}_{\hbar}(\mathrm{GL}_{N})\cong\mathrm{U}_{\hbar}(\mathfrak{gl}_{N}^{R})\otimes\mathcal{O}(\mathrm{GL}_{N})$ is a free module, it follows from trivializations (4.8) and (5.3) that this map is injective. By exactness of the Kostant-Whittaker reduction [BF08, Lemma 4(a)], the induced map
@@
The following theorem is an explicit presentation of the isomorphism from [BFN19, Appendix B].
@@
${\mathrm{Y}^{0}_{-2N}(\mathfrak{sl}_{2})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\sim}$$\scriptstyle{\mathrm{GKLO}_{N}}$${N^{L}_{+}\,\backslash\!\!\backslash_{-\psi}\,\mathrm{D}_{\hbar}(\mathrm{GL}_{N})/\!\!/_{\psi}\,N^{R}_{+}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathrm{Miura}_{N}}$${{}^{N}\tilde{\mathcal{A}}_{\hbar}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\sim}$${N^{L}_{+}\,\backslash\!\!\backslash_{-\psi}\,\mathrm{D}_{\hbar}^{\mathrm{gen}}(\mathrm{GL}_{N})/\!\!/N^{R}_{-}}$
@@
where the first map is (6.4). By naturality of the Miura transform and Theorem˜5.9, the square
${\mathrm{Y}_{-N,N}(\mathfrak{gl}_{2})^{\geq 0}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathrm{GKLO}_{N}}$${N^{L}_{+}\,\backslash\!\!\backslash_{-\psi}\,\mathrm{D}_{\hbar}(\mathrm{GL}_{N})/\!\!/_{\psi}\,N^{R}_{+}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathrm{Miura}_{N}}$${{}^{N}\tilde{\mathcal{A}}_{\hbar}^{-}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${N^{L}_{+}\,\backslash\!\!\backslash_{-\psi}\,\mathrm{D}_{\hbar}^{\mathrm{gen}}(\mathrm{GL}_{N})/\!\!/N^{R}_{-}}$
@@
${\mathrm{Y}_{-N,N}(\mathfrak{gl}_{2})^{\leq 0}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathrm{GKLO}_{N}}$${N^{L}_{+}\,\backslash\!\!\backslash_{-\psi}\,\mathrm{D}_{\hbar}(\mathrm{GL}_{N})/\!\!/_{\psi}\,N^{R}_{+}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathrm{Miura}_{N}}$${{}^{N}\tilde{\mathcal{A}}_{\hbar}^{+}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${N^{L}_{+}\,\backslash\!\!\backslash_{-\psi}\,\mathrm{D}_{\hbar}^{\mathrm{gen}}(\mathrm{GL}_{N})/\!\!/N^{R}_{-}}$
by Theorem˜B.3. To check the remaining relation (2.9), we can use ˜7.8 and check it inside
@@
which is a straightforward calculation, for instance, see [BFN19, Appendix B(v)]. Therefore, we have a map
@@
whose composition with the Miura transform (7.3) coincides with the GKLO map. It is surjective by the first part of ˜7.5. By ˜7.8, the target is identified with its image in $^{N}\tilde{\mathcal{A}}_{\hbar}$. Then ˜2.17 gives an isomorphism
@@
In view of Remark˜6.5, the $W$-action on $N^{L}_{+}\,\backslash\!\!\backslash_{-\psi}\,\mathrm{D}_{\hbar}(\mathrm{GL}_{N})/\!\!/N_{-}\cong{}^{N}\tilde{\mathcal{A}}_{\hbar}$ is